# Stabilized symplectic embeddings of higher-dimensional ellipsoids

### v1

@@ 1 Introduction and statements of results
A symplectic embedding problem asks whether one symplectic manifold can be symplectically embedded into another. Beyond their intrinsic interest, symplectic embeddings can be used to define new global invariants for symplectic manifolds [MR3777016, MR2369441]. Due to the co-existence of rigidity and flexibility in symplectic embeddings, studying symplectic embedding problems helps to illustrate the boundary between rigidity and flexibility in symplectic geometry. Symplectic embedding techniques have also found intriguing applications to classical dynamical systems, such as the restricted three-body problem [Frauenfelder_2018]; see also [MR3777016, Section 4.5].
@@
Problem 1.1 has been completely solved in dimension four [McDuff-Hoferconjecture, McDuff-Schlenk, Faisal:2024aa]. However, little is known in higher dimensions; we recommend [Buse-Hindellipsoidsindimensiongreaterthanfour, Cristofaro-Hind-Siegel-stabilizedembedding, MR3777016, Schlenkellipsoids] for some progress and open problems.
@@
for any $1\leq x_{2}\leq x_{3}\leq\dots\leq x_{n}\leq\infty$, where $\xrightarrow[]{s}$” represents a symplectic embedding with the standard symplectic structures on both the domain and target. We mention that function (1.2), for $k=2$, has already appeared in the work of Hind [MR3394319, Section 1.2].
@@
  • McDuff–Schlenk [McDuff-Schlenk]: the function $\mathcal{EB}_{2}(\cdot)$ is understood completely. On the interval $0,\tau^{4}]$, where $\tau=\frac{1+\sqrt{5}}{2}$, the function $\mathcal{EB}_{2}(x_{2})$ is given by the Fibonacci stairs. For any $x_{2}\in[\tau^{4},7]$, we have $3\mathcal{EB}_{2}(x_{2})=(x_{2}+1)$. For all $x_{2}\geq 8\frac{1}{36}$, we have $\mathcal{EB}_{2}(x_{2})=\sqrt{x_{2}}$. On the interval $7,8\frac{1}{36}]$, the function $\mathcal{EB}_{2}(x_{2})$ takes a complicated form; see [McDuff-Schlenk, Theorem 1.1.2].
  • Buse–Hind [Buse-Hindellipsoidsindimensiongreaterthanfour]: for six-dimensional ellipsoids $E^{6}(1,x_{2},x_{3})$ satisfying $x_{2}^{2}+x_{3}^{2}\geq 1.41\times 10^{101}$, the volume constraint is optimal: $\mathcal{EB}_{3}(x_{2},x_{3})=\sqrt[3]{x_{2}x_{3}}$. Also for small values of $x_{2}$ and $x_{3}$, the values $\mathcal{EB}_{3}(x_{2},x_{3})$ are known; see [Buse-Hindellipsoidsindimensiongreaterthanfour, Lemma 2.7—2.11, Figure 1].
  • Cristofaro–Hind–McDuff [MR3789827]: on the interval $0,\tau^{4}]$, where $\tau=\frac{1+\sqrt{5}}{2}$, the function $\mathcal{EC}_{n-2}(x_{2},\infty,\dots,\infty)$ is given by the Fibonacci stairs for any $n\geq 3$.
  • Hind [MR3394319]: For $x_{2}>\tau^{4}$ and any $n\geq 3$ it holds that
@@
  • Hind–Kerman [Hind-Kerman-ellipembedding, MR3867635]: for $x_{2}=3l-1$, where $l$ is an odd indexed Fibonacci number, we have
@@
  • McDuff [MR3782228]: for $x_{2}=3l-1$, where $l$ is any positive integer, we have
@@
  • McDuff–Siegel [McDuff:2024aa]: for any integer $n\geq 3$ and all $x_{2}\geq\tau^{4}$, we have
@@ 1.1 Main results
Inspired by Gromov’s pioneering idea in the proof of his non-squeezing theorem, one can seek sharp numerical obstructions to symplectic embeddings by constructing pseudo-holomorphic curves with carefully controlled symplectic area. In case of stabilized embeddings of ellipsoids, this approach leads to the problem of producing pseudo-holomorphic curves of arbitrarily large degree, subject to specific geometric and asymptotic constraints, in complex projective spaces; see [Hind-Kerman-ellipembedding, McDuff:2024aa, Mikhalkin:2023aa, MR4466005, McDuff:2023aa] and the references therein. In particular, a proof of (1.5) reduces to establishing the existence of certain curves of arbitrarily large degree and a special type of singularities [MR3782228, McDuff:2024ab, MR4466005, McDuff:2024aa].
@@
By [Pereira:2025aa, Theorem 4.37], the Lagrangian capacity of an ellipsoid is given by
@@
By the monotonicity property of the Lagrangian capacity [Cieliebak2018, Section 1.2], for any positive integers $n$ and $k$ such that $n-k\geq 1$, we have
@@
for all $c\leq x_{2}\leq x_{3}\leq\dots\leq x_{n}$. However, by Buse–Hind [Buse-Hindellipsoidsindimensiongreaterthanfour, Theorem 2.1], the obstruction to embeddings of large six-dimensional ellipsoids into a six-dimensional ball coming from volumes is optimal, i.e., $\mathcal{EB}_{3}(x_{2},x_{3})=\sqrt[3]{x_{2}x_{3}}$ for large $x_{2},x_{3}$. In particular, this means that the lower bound in (1.9) is far from being sharp and is much weaker than the lower bound coming from the volume constraint.
Theorem 1.2 does not cover the case $k=2$. To address this remaining case, we adopt a different strategy. The key observation is that a symplectic embedding problem can often be reformulated either as a Lagrangian embedding problem or as a classification problem for monotone Lagrangians. This perspective offers two advantages. First, it enables the use of algebraic techniques developed for monotone Lagrangians to study symplectic embedding questions. Second, it allows one to apply the splitting argument of Cieliebak–Mohnke [Cieliebak2018], which may yield sharp symplectic embedding obstructions by employing low-degree holomorphic curves with local tangency constraints. Following this, we reformulate the stabilized symplectic embedding problem for ellipsoids as a classification problem for monotone Lagrangian tori in complex projective spaces. More precisely, we offer the following to determine the asymptotic of (1.2) using our approach.
@@ 2 Holomorphic planes with ends on a skinny ellipsoid
We start with a review of some fundamental computations for standard ellipsoids. Our main reference is [MR3868228, Section 2.1].
@@
Theorem 2.2 (Gutt–Hutchings [MR3868228, Section 2.1]).
@@
This notion of tangency does not depend on the choice of the functions $h$ and $g$. It only depends on the germ of $D$ near $p$; see [Cieliebak_2007, Section $6$] for a proof.
By [Cieliebak_2007, Lemma 7.1], the tangency constraint $\ll\mathcal{T}_{D}^{k-1}p\gg$ can be interpreted as a local intersection number of the image of $u$ with the divisor $D$ as follows: Choose [^1] The symplectic form $\omega$ is exact on $O(p)$, so $u$ cannot be contained in the divisor $D$ by Stokes’ theorem. a small ball $B$ around $z_{0}$ such that $u^{-1}(D)\cap B=\{z_{0}\}$. Smoothly perturb $u|_{B}$ away from $\partial B$ to make it transverse to $D$. The signed count of transverse intersections of $u|_{B}$ with $D$ equals $k$.
@@
Theorem 2.6 ([Pereira:2025aa, Section 4.6], cf. [MR4332489, Lemma 4.1.3, Lemma 3.1.11], [Mikhalkin:2023aa, Theorem 5.2.1]).
@@
    By [Faisal:2024, Corollary 3.10], for each $J_{j}$, there exists a $J_{j}$-holomorphic sphere in $\mathbb{CP}^{k}\times T^{2(n-k)}$ in the homology class $\mathbb{CP}^{1}\times\{*\}]$ satisfying the constraint $\ll\mathcal{T}_{D_{1}}^{k-1}0\gg$, where the local symplectic divisor $D_{1}$ is defined by the equation $z_{1}=0$ in the embedded ellipsoid $E^{2n}(\epsilon\vec{x})$.
@@
    We start by proving that the holomorphic building $\mathbb{H}$ has no node between its non-constant curve components. In particular, the building does not have a node at the constrained marked point $0\in E^{2n}(\epsilon\vec{x})$. Since $\mathbb{H}$ has genus zero, any node between non-constant components decomposes the building $\mathbb{H}$ into two pieces $A_{1}$ and $A_{2}$. These are represented by non-constant holomorphic curves (lying possibly at different levels). By [Cieliebak_2005, Lemma 2.6], we have $\Omega(A_{1})>0,\Omega(A_{2})>0$, where $\Omega=\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{std}}$. The building $\mathbb{H}$ has exactly one intersection with the $J_{\infty}$-holomorphic hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$ because of the positivity of intersection and the fact that $\mathbb{H}$ is the result of breaking of a holomorphic sphere that has intersection number +1+1 with $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. Therefore, one of these spheres, say $A_{1}$, lies in the complement of $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. But since $\Omega=\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{std}}$ is exact in the complement of the hypersuface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$, we must have $\Omega(A_{1})=0$ by Stokes’ theorem. This is a contradiction. We conclude that a smooth component, denoted by $u_{\mathrm{bot}}$, in the bottom level $\widehat{E^{2n}}(\epsilon\vec{x})$ inherits the tangency constraint $\ll\mathcal{T}_{D_{1}}^{k-1}0\gg$.
@@
    Recall from Step 04 that the bottom level $\widehat{E^{2n}}(\epsilon\vec{x})$ contains a smooth component, denoted by $u_{\mathrm{bot}}$, that inherits the tangency constraint $\ll\mathcal{T}_{D_{1}}^{k-1}0\gg$. We show that $u_{\mathrm{bot}}$ has a single positive puncture that is asymptotic to a short Reeb orbit, i.e., to a cover of $\beta_{1}$.
@@
    On the other hand, gluing $u_{\infty}$ and $u_{\mathrm{bot}}$ produces a curve that computes the curve count in [Faisal:2024, Corollary 3.10]. Therefore, by [Faisal:2024, Corollary 3.10], we have
@@ 3 Enumerative descendants
Let $(X,\omega)$ be a compact monotone symplectic manifold, and let $d$ be a positive integer. Following [Cieliebak_2007, Cieliebak2018], we introduce a variant of Gromov–Witten invariants that we denote by $\langle\psi_{d-2}p\rangle_{X,d}^{\bullet}$ and refer to as enumerative descendants of $(X,\omega)$. The definition goes as follows.
@@
The coherent perturbation $K$ (cf. [Cieliebak2018, Section 6.3–4, and p. 274]) in the holomorphic curve equation is given by $K=X_{H}\otimes\beta$ , where
@@
By [Cieliebak_2007, Theorem 1.2–3] or [Tonkonog:2018aa, Section 2.3], the count defined by (3.2) does not depend on $p,D,J$, $\Sigma$, and the coherent perturbation $K$.
By [Faisal:2024, Corollary 3.10], for $(X,\omega)=(\mathbb{CP}^{k}\times T^{2m},\omega_{\mathrm{FS}}\oplus\mathbb{\omega}_{\mathrm{std}})$ and the indecomposable homology class $A=[\mathbb{CP}^{1}\times\{*\}]\in H_{2}(\mathbb{CP}^{k}\times T^{2m},\mathbb{Z})$, the count $\langle\psi_{k-1}p\rangle_{\mathbb{CP}^{k}\times T^{2m},k+1}^{\bullet}$ is equal to (k1)!(k-1)! when $K=0$, so we have the following.
Theorem 3.1 ([Faisal:2024, Corollary 3.10], [Cieliebak_2007, Theorem 1.2–3]).
@@ 4 Gravitational descendants of cotangent bundles
Let $(X,\lambda)$ be a non-degenerate Liouville domain. We denoted by $\operatorname{SH^{*}_{S^{1},+}}$ the the positive $S^{1}$-equivariant symplectic cohomology of $(X,\lambda)$. We briefly recall its definition following [MR3671507]. Let
@@
To achieve $\partial\circ\partial=0$ and make the chain complex $\operatorname{CF^{*}_{S^{1},+}}(X)$ independent (up to chain homotopy) on the choice of the almost complex structure $J$, one requires a suitable virtual perturbation scheme to define the curve count involved; for a proposal see [Chaidez:2024aa, Pardon-Contacthomologyandvirtualfundamentalcycles]. When $(X,\lambda)$ is the unit codisk bundle of a closed Riemannian manifold that admits a metric of non-positive sectional curvature, then there are no closed Reeb orbits on $\partial X$ that are contractible in $X$, as in this case, there are no contractible closed geodesics. The differential $\partial$ counts pure holomorphic cylinders, which are unbranched by the Riemann–Hurwitz formula, interpolating between the input and output closed Reeb orbits. In this case, one does not require any virtual perturbation scheme to define $\operatorname{CF^{*}_{S^{1},+}}(X)$; in particular, this is the case for the unit codisk bundle $D^{*}T^{n}$ of the torus $T^{n}$ [MR2475400, MR3671507] in which we are interested in this document.
@@
From now on, we assume that $(X,\lambda)$ is the unit codisk bundle $D^{*}L$ of a closed Riemannian manifold $L$ of dimension $n$ that admits a metric of non-positive sectional curvature. Up to an arbitrary high length (action) truncation, by [Cieliebak2018, Lemma 2.2], we assume that every closed geodesic (closed Reeb orbit) has Morse index (Conley–Zehnder index) between 0 and n1n-1. From now on, we assume that all the generators of the chain complex $\operatorname{CF^{*}_{S^{1},+}}(D^{*}L)$ have actions smaller than a fixed number $A>0$. The grading on $\operatorname{CF^{*}_{S^{1},+}}(D^{*}L)$ used in this document is the one used in [Tonkonog:2018aa]. This is given by
@@
Here, the perturbation $X_{H}\otimes\beta$ in the holomorphic curve equation is chosen as in Section 3. These moduli spaces are generically transversally cut out and are, moreover, rigid. Following [Tonkonog:2018aa, Section 4.4], these rigid moduli spaces yield linear maps
@@
whenever $\gamma_{i}$ are generators of $\operatorname{CF^{0}_{S^{1},+}}(D^{*}L)$ and extend linearly to the full complex. In fact, by [Tonkonog:2018aa, Theorem 4.2], the maps $\langle\cdot|\cdots|\cdot\rangle$ belong to a $2$-family of linear maps $\{\psi_{m-1}^{k}\}_{m\geq 1,k\geq 2}$ that defines, for each integer $m\geq 1$, an $L_{\infty}$-algebra structure on the full complex $\operatorname{CF^{*}_{S^{1},+}}(X)$, where $X$ belongs to a more general class of Liouville domains. These linear operations are called gravitational descendants in [Tonkonog:2018aa].
By [Tonkonog:2018aa, Proposition 4.4], for each $k\geq 2$ the maps $\langle\cdot|\cdots|\cdot\rangle:\operatorname{CF^{0}_{S^{1},+}}(D^{*}L)^{\otimes k}\to\mathbb{Z}$ descend to the cohomological level operations
@@
Theorem 4.1 (Descendants of the torus $T^{n}$, [Tonkonog:2018aa, Theorem 4.5]).
@@
That is, the existence of the embedding (5.4) satisfying (5.5) implies the existence of too few curves and hence a contradiction. To prove the latter claim, we follow the structure of arguments from [Faisal:2024, Section 6].
@@
    Take a flat metric on $L_{\Phi_{i},r_{i}}$. After scaling it, we can symplectically embed the codisk bundle of radius $2$, denoted by $D^{*}_{2}L_{\Phi_{i},r_{i}}$, into $\mathbb{CP}^{k}\times T^{2(n-k)}$ in the complement of the hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. Perturb this metric according to [Cieliebak2018, Lemma 2.2] to a Riemannian metric $g$ such that, with respect to $g$, every closed geodesic $\gamma$ of length less than or equal to $c=r_{i}$ is noncontractible, nondegenerate (as a critical point of the energy functional) and satisfies
@@
    Following the arguments of [Faisal:2024, Section 6.3] for the monotone torus $L_{\Phi_{i},r_{i}}$, we conclude that

### v2

@@ 1 Introduction and statements of results
A symplectic embedding problem asks whether one symplectic manifold can be symplectically embedded into another. Beyond their intrinsic interest, symplectic embeddings can be used to define new global invariants for symplectic manifolds [SCH18, CHL+07]. Due to the co-existence of rigidity and flexibility in symplectic embeddings, studying symplectic embedding problems helps to illustrate the boundary between rigidity and flexibility in symplectic geometry. Symplectic embedding techniques have also found intriguing applications to classical dynamical systems, such as the restricted three-body problem [Fv18]; see also [SCH18, Section 4.5].
@@
Problem 1.1 has been completely solved in dimension four [MCD11, MS12, 12]. However, little is known in higher dimensions; we recommend [BH11, CHS22, SCH18, SCH03] for some progress and open problems.
@@
for any $1\leq x_{2}\leq x_{3}\leq\dots\leq x_{n}\leq\infty$, where $\xrightarrow[]{s}$” represents a symplectic embedding with the standard symplectic structures on both the domain and target. We mention that function (1.2), for $k=2$, has already appeared in the work of Hind [HIN15, Section 1.2].
@@
  • McDuff–Schlenk [MS12]: the function $\mathcal{EB}_{2}(\cdot)$ is understood completely. On the interval $0,\tau^{4}]$, where $\tau=\frac{1+\sqrt{5}}{2}$, the function $\mathcal{EB}_{2}(x_{2})$ is given by the Fibonacci stairs. For any $x_{2}\in[\tau^{4},7]$, we have $3\mathcal{EB}_{2}(x_{2})=(x_{2}+1)$. For all $x_{2}\geq 8\frac{1}{36}$, we have $\mathcal{EB}_{2}(x_{2})=\sqrt{x_{2}}$. On the interval $7,8\frac{1}{36}]$, the function $\mathcal{EB}_{2}(x_{2})$ takes a complicated form; see [MS12, Theorem 1.1.2].
  • Buse–Hind [BH11]: for six-dimensional ellipsoids $E^{6}(1,x_{2},x_{3})$ satisfying $x_{2}^{2}+x_{3}^{2}\geq 1.41\times 10^{101}$, the volume constraint is optimal: $\mathcal{EB}_{3}(x_{2},x_{3})=\sqrt[3]{x_{2}x_{3}}$. Also for small values of $x_{2}$ and $x_{3}$, the values $\mathcal{EB}_{3}(x_{2},x_{3})$ are known; see [BH11, Lemma 2.7—2.11, Figure 1].
  • Cristofaro–Hind–McDuff [CHM18]: on the interval $0,\tau^{4}]$, where $\tau=\frac{1+\sqrt{5}}{2}$, the function $\mathcal{EC}_{n-2}(x_{2},\infty,\dots,\infty)$ is given by the Fibonacci stairs for any $n\geq 3$.
  • Hind [HIN15]: For $x_{2}>\tau^{4}$ and any $n\geq 3$ it holds that
@@
  • Hind–Kerman [HK14, HK18]: for $x_{2}=3l-1$, where $l$ is an odd indexed Fibonacci number, we have
@@
  • McDuff [MCD18]: for $x_{2}=3l-1$, where $l$ is any positive integer, we have
@@
  • McDuff–Siegel [MS24b]: for any integer $n\geq 3$ and all $x_{2}\geq\tau^{4}$, we have
@@ 1.1 Main results
Inspired by Gromov’s pioneering idea in the proof of his non-squeezing theorem, one can seek sharp numerical obstructions to symplectic embeddings by constructing pseudo-holomorphic curves with carefully controlled symplectic area. In case of stabilized embeddings of ellipsoids, this approach leads to the problem of producing pseudo-holomorphic curves of arbitrarily large degree, subject to specific geometric and asymptotic constraints, in complex projective spaces; see [HK14, MS24b, MS23, SIE22, MS23] and the references therein. In particular, a proof of (1.5) reduces to establishing the existence of certain curves of arbitrarily large degree carrying a special type of singularities [MCD18, MS24a, SIE22, MS24b].
@@
By [PER25, Theorem 4.37], the Lagrangian capacity of an ellipsoid is given by
@@
By the monotonicity property of the Lagrangian capacity [CM18, Section 1.2], for any positive integers $n$ and $k$ such that $n-k\geq 1$, we have
@@
for all $c\leq x_{2}\leq x_{3}\leq\dots\leq x_{n}$. However, by Buse–Hind [BH11, Theorem 2.1], the obstruction to embeddings of large six-dimensional ellipsoids into a six-dimensional ball coming from volumes is optimal, i.e., $\mathcal{EB}_{3}(x_{2},x_{3})=\sqrt[3]{x_{2}x_{3}}$ for large $x_{2},x_{3}$. In particular, this means that the lower bound in (1.9) is far from being sharp and is much weaker than the lower bound coming from the volume constraint.
Theorem 1.2 does not cover the case $k=2$. To address this remaining case, we adopt a different strategy. The key observation is that a symplectic embedding problem can often be reformulated either as a Lagrangian embedding problem or as a classification problem for monotone Lagrangians. This perspective offers two advantages. First, it enables the use of algebraic techniques developed for monotone Lagrangians to study symplectic embedding questions. Second, it allows one to apply the splitting argument of Cieliebak–Mohnke [CM18], which may yield sharp symplectic embedding obstructions by employing low-degree holomorphic curves with local tangency constraints. Following this, we reformulate the stabilized symplectic embedding problem for ellipsoids as a classification problem for monotone Lagrangian tori in complex projective spaces. More precisely, we offer the following to determine the asymptotic of (1.2) using our approach.
@@ 2 Holomorphic planes with ends on a skinny ellipsoid
We start with a review of some fundamental computations for standard ellipsoids. Our main reference is [GH18, Section 2.1].
@@
Theorem 2.2 (Gutt–Hutchings [GH18, Section 2.1]).
@@
This notion of tangency does not depend on the choice of the functions $h$ and $g$. It only depends on the germ of $D$ near $p$; see [CM07, Section $6$] for a proof.
By [CM07, Lemma 7.1], the tangency constraint $\ll\mathcal{T}_{D}^{k-1}p\gg$ can be interpreted as a local intersection number of the image of $u$ with the divisor $D$ as follows: Choose [^1] The symplectic form $\omega$ is exact on $O(p)$, so $u$ cannot be contained in the divisor $D$ by Stokes’ theorem. a small ball $B$ around $z_{0}$ such that $u^{-1}(D)\cap B=\{z_{0}\}$. Smoothly perturb $u|_{B}$ away from $\partial B$ to make it transverse to $D$. The signed count of transverse intersections of $u|_{B}$ with $D$ equals $k$.
@@
Theorem 2.6 ([PER25, Section 4.6], cf. [MS21, Lemma 4.1.3, Lemma 3.1.11], [MS23, Theorem 5.2.1]).
@@
    By [FAI25, Corollary 3.10], for each $J_{j}$, there exists a $J_{j}$-holomorphic sphere in $\mathbb{CP}^{k}\times T^{2(n-k)}$ in the homology class $\mathbb{CP}^{1}\times\{*\}]$ satisfying the constraint $\ll\mathcal{T}_{D_{1}}^{k-1}0\gg$, where the local symplectic divisor $D_{1}$ is defined by the equation $z_{1}=0$ in the embedded ellipsoid $E^{2n}(\epsilon\vec{x})$.
@@
    We start by proving that the holomorphic building $\mathbb{H}$ has no node between its non-constant curve components. In particular, the building does not have a node at the constrained marked point $0\in E^{2n}(\epsilon\vec{x})$. Since $\mathbb{H}$ has genus zero, any node between non-constant components decomposes the building $\mathbb{H}$ into two pieces $A_{1}$ and $A_{2}$. These are represented by non-constant holomorphic curves (lying possibly at different levels). By [CM05, Lemma 2.6], we have $\Omega(A_{1})>0,\Omega(A_{2})>0$, where $\Omega=\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{std}}$. The building $\mathbb{H}$ has exactly one intersection with the $J_{\infty}$-holomorphic hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$ because of the positivity of intersection and the fact that $\mathbb{H}$ is the result of breaking of a holomorphic sphere that has intersection number +1+1 with $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. Therefore, one of these spheres, say $A_{1}$, lies in the complement of $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. But since $\Omega=\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{std}}$ is exact in the complement of the hypersuface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$, we must have $\Omega(A_{1})=0$ by Stokes’ theorem. This is a contradiction. We conclude that a smooth component, denoted by $u_{\mathrm{bot}}$, in the bottom level $\widehat{E^{2n}}(\epsilon\vec{x})$ inherits the tangency constraint $\ll\mathcal{T}_{D_{1}}^{k-1}0\gg$.
@@
    Recall from Step 04 that the bottom level $\widehat{E(a,b)}(\epsilon\vec{x})$ contains a smooth component, denoted by $u_{\mathrm{bot}}$, that inherits the tangency constraint $\ll\mathcal{T}_{D_{1}}^{k-1}0\gg$. We show that $u_{\mathrm{bot}}$ has a single positive puncture that is asymptotic to a short Reeb orbit, i.e., to a cover of $\beta_{1}$.
@@
    On the other hand, gluing $u_{\infty}$ and $u_{\mathrm{bot}}$ produces a curve that computes the curve count in [FAI25, Corollary 3.10]. Therefore, by [FAI25, Corollary 3.10], we have
@@ 3 Enumerative descendants
Let $(X,\omega)$ be a compact monotone symplectic manifold, and let $d$ be a positive integer. Following [CM07, CM18], we introduce a variant of Gromov–Witten invariants that we denote by $\langle\psi_{d-2}p\rangle_{X,d}^{\bullet}$ and refer to as enumerative descendants of $(X,\omega)$. The definition goes as follows.
@@
The coherent perturbation $K$ (cf. [CM18, Section 6.3–4, and p. 274]) in the holomorphic curve equation is given by $K=X_{H}\otimes\beta$ , where
@@
By [CM07, Theorem 1.2–3] or [TON18, Section 2.3], the count defined by (3.2) does not depend on $p,D,J$, $\Sigma$, and the coherent perturbation $K$.
By [FAI25, Corollary 3.10], for $(X,\omega)=(\mathbb{CP}^{k}\times T^{2m},\omega_{\mathrm{FS}}\oplus\mathbb{\omega}_{\mathrm{std}})$ and the indecomposable homology class $A=[\mathbb{CP}^{1}\times\{*\}]\in H_{2}(\mathbb{CP}^{k}\times T^{2m},\mathbb{Z})$, the count $\langle\psi_{k-1}p\rangle_{\mathbb{CP}^{k}\times T^{2m},k+1}^{\bullet}$ is equal to (k1)!(k-1)! when $K=0$, so we have the following.
Theorem 3.1 ([FAI25, Corollary 3.10], [CM07, Theorem 1.2–3]).
@@ 4 Gravitational descendants of cotangent bundles
Let $(X,\lambda)$ be a non-degenerate Liouville domain. We denoted by $\operatorname{SH^{*}_{S^{1},+}}$ the the positive $S^{1}$-equivariant symplectic cohomology of $(X,\lambda)$. We briefly recall its definition following [BO17]. Let
@@
To achieve $\partial\circ\partial=0$ and make the chain complex $\operatorname{CF^{*}_{S^{1},+}}(X)$ independent (up to chain homotopy) on the choice of the almost complex structure $J$, one requires a suitable virtual perturbation scheme to define the curve count involved; for a proposal see [CHA25, PAR19]. When $(X,\lambda)$ is the unit codisk bundle of a closed Riemannian manifold that admits a metric of non-positive sectional curvature, then there are no closed Reeb orbits on $\partial X$ that are contractible in $X$, as in this case, there are no contractible closed geodesics. The differential $\partial$ counts pure holomorphic cylinders, which are unbranched by the Riemann–Hurwitz formula, interpolating between the input and output closed Reeb orbits. In this case, one does not require any virtual perturbation scheme to define $\operatorname{CF^{*}_{S^{1},+}}(X)$; in particular, this is the case for the unit codisk bundle $D^{*}T^{n}$ of the torus $T^{n}$ [BO09, BO17] in which we are interested in this document.
@@
From now on, we assume that $(X,\lambda)$ is the unit codisk bundle $D^{*}L$ of a closed Riemannian manifold $L$ of dimension $n$ that admits a metric of non-positive sectional curvature. Up to an arbitrary high length (action) truncation, by [CM18, Lemma 2.2], we assume that every closed geodesic (closed Reeb orbit) has Morse index (Conley–Zehnder index) between 0 and n1n-1. From now on, we assume that all the generators of the chain complex $\operatorname{CF^{*}_{S^{1},+}}(D^{*}L)$ have actions smaller than a fixed number $A>0$. The grading on $\operatorname{CF^{*}_{S^{1},+}}(D^{*}L)$ used in this document is the one used in [TON18]. This is given by
@@
Here, the perturbation $X_{H}\otimes\beta$ in the holomorphic curve equation is chosen as in Section 3. These moduli spaces are generically transversally cut out and are, moreover, rigid. Following [TON18, Section 4.4], these rigid moduli spaces yield linear maps
@@
whenever $\gamma_{i}$ are generators of $\operatorname{CF^{0}_{S^{1},+}}(D^{*}L)$ and extend linearly to the full complex. In fact, by [TON18, Theorem 4.2], the maps $\langle\cdot|\cdots|\cdot\rangle$ belong to a $2$-family of linear maps $\{\psi_{m-1}^{k}\}_{m\geq 1,k\geq 2}$ that defines, for each integer $m\geq 1$, an $L_{\infty}$-algebra structure on the full complex $\operatorname{CF^{*}_{S^{1},+}}(X)$, where $X$ belongs to a more general class of Liouville domains. These linear operations are called gravitational descendants in [TON18].
By [TON18, Proposition 4.4], for each $k\geq 2$ the maps $\langle\cdot|\cdots|\cdot\rangle:\operatorname{CF^{0}_{S^{1},+}}(D^{*}L)^{\otimes k}\to\mathbb{Z}$ descend to the cohomological level operations
@@
Theorem 4.1 (Descendants of the torus $T^{n}$, [TON18, Theorem 4.5]).
@@
That is, the existence of the embedding (5.4) satisfying (5.5) implies the existence of too few curves and hence a contradiction. To prove the latter claim, we follow the structure of arguments from [FAI25, Section 6].
@@
    Take a flat metric on $L_{\Phi_{i},r_{i}}$. After scaling it, we can symplectically embed the codisk bundle of radius $2$, denoted by $D^{*}_{2}L_{\Phi_{i},r_{i}}$, into $\mathbb{CP}^{k}\times T^{2(n-k)}$ in the complement of the hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. Perturb this metric according to [CM18, Lemma 2.2] to a Riemannian metric $g$ such that, with respect to $g$, every closed geodesic $\gamma$ of length less than or equal to $c=r_{i}$ is noncontractible, nondegenerate (as a critical point of the energy functional) and satisfies
@@
    Following the arguments of [FAI25, Section 6.3] for the monotone torus $L_{\Phi_{i},r_{i}}$, we conclude that